OpenStudy (anonymous):

What is the solution to the rational equation 3 over 2 - 1 over x plus 5 = negative 2 over 2x plus 10

OpenStudy (anonymous):

\[\frac{3}{2} -\frac{1}{x}+5=\frac{-2}{2x} +10\] is this your problem?

OpenStudy (anonymous):

1/x+5 and -2 /2x+10